OpenStudy (anonymous):

Deepak bought a new car at the dealership for $25,000. It is estimated that the value of the car will decrease 9% each year. Which exponential function models the value v of the car after t years?

OpenStudy (anonymous):

@Hero @JackJordan

hero (hero):

Says "which" meaning, a list of answer choices are available to choose from.

OpenStudy (anonymous):

a) v=25,000(1.9)^t b) (same^^) (0.1)^t c) (same^^) (0.91)^t d) (same^^) (1.09)^t

OpenStudy (anonymous):

i know its either a or d because the 1 comes before the decimal, but i got confused because i don't remember if the 9 has a 0 infront or not

hero (hero):

It's not proper to post answer choices in that manner.

hero (hero):

How do you figure that it's either a or d?

hero (hero):

Explain how you eliminated b or c.

OpenStudy (anonymous):

a) v=25,000(1.9)^t b) v=25,000(0.1)^t c) v=25,000(0.91)^t d) v=25,000(1.09)^t my b

OpenStudy (anonymous):

i remember that the 1 comes before the decimal

hero (hero):

1 comes before the decimal for exponential INCREASE.

OpenStudy (anonymous):

oh. so when it's not increasing, the 1 comes after the decimal?

OpenStudy (anonymous):

? @Hero

hero (hero):

Basically, the only the only value of interest for the general formula \(v = 25000^{kt}\) is the k value. If there's an exponential increase, then \(k = 1 + r\) . If there's an exponential decrease, then \(k = 1 - r\) In this case \(r = 0.09\).

OpenStudy (anonymous):

so it's (0.91)^t

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@mathstudent55

hero (hero):

The general formula is \(v = 25000k^t\) sorry.

OpenStudy (mathstudent55):

The formula for exponential growth or decay is: F = P(1 + r)^t where F = future amount P = present amount r = rate of growth or decay t = number of years If r is a rate of growth, r is positive. If r is a rate of decay, r is negative. You have a rate of decay, so r = -9% = -0.09 \(F = 25,000(1 - 0.9)^t\) \(F = 25,000(0.91)^t\)

OpenStudy (anonymous):

perfect! thank you!

OpenStudy (mathstudent55):

Yw

hero (hero):

I wrote the k as an exponent. Sorry about that.